CEPS Group just paid an annual dividend of OMR 3.55 a share. The firm plans to increase the annual dividends by 3 percent over the next 3 years. After that time, the dividends will be held constant at OMR 4.5 per share. What is this stock worth today at a 7 percent discount rate?
Answer #1

Year 1 dividend = 3.55 (1 + 3%) = 3.6565

Year 2 dividend = 3.6565 (1 + 3%) = 3.766195

Year 3 dividend = 3.766195 (1 + 3%) = 3.879181

Year 4 dividend = 4.5

Value at year 3 = D4 / required rate

Value at year 3 = 4.5 / 0.07

Value at year 3 = 64.285714

Stock price = Present value of cash inflows

Stock price = 3.6565 / (1 + 0.07)^1 + 3.766195 / (1 + 0.07)^2 + 3.879181 / (1 + 0.07)^3 + 64.285714 / (1 + 0.07)^3

Stock price = $62.35